OpenStudy (anonymous):

36x2 − 12x divided by 4x

OpenStudy (anonymous):

You would get better help in the Mathematics section however, 36x2/4x=9x2 12x/4x=3x 9x2-3x=?

OpenStudy (kmazraee):

15

OpenStudy (anonymous):

lol no 9x2 and 3x are factors, it was not 9 • 2 - 3

OpenStudy (shadowfiend):

It's not just a matter of getting better help. @dc333, please don't post questions that aren't relevant to the subject you're in. In this case, if you want to ask a math question, please post it in the math group at http://openstudy.com/groups/mathematics . @kmarazee, please don't provide answers without explanations, even if they're wrong. Rebecca was trying to explain how to arrive at the solution. Please follow this example. See our Code of Conduct at http://openstudy.com/code-of-conduct for more.

OpenStudy (kmazraee):

of course sir! thanks a lot!